events c and d are depentdable events.
p(c)=0.75
p(d given c)=0.5
p(c and d)=?

Answer by robertb(5830)

c and being dependent events,
P(c AND d) = P(d GIVEN c)*P(c) = 0.5*0.75 = 0.375